The average number of calories in a 1.5-ounce chocolate bar is 217. Suppose that the distribution of calories is approximately normal with standard deviation is 9. Find the probability that a randomly selected chocolate bar will have more than 190.




1
Expert's answer
2022-06-13T15:30:22-0400

"P(X>190)=P(Z>\frac{190-217}{9})=P(Z>-3)=P(Z<3)=0.9987."